# Information Sifting Funnel: Privacy-preserving Collaborative Inference Against Model Inversion Attacks

### v1

@@ 6.3. Experimental Results
1) The effect of various defenses on the MI and $\delta(z)$. the effectiveness of defenses is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$ and the theoretical analysis presented in Section 4. Specifically, unprotected edge models show significant vulnerability to input reconstruction under both types of MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98. When protected by Perturbation-based defenses, all methods except NLS exhibit over a 10% drop in Test ACC. Moreover, these methods achieve only an 11% reduction in MI at most, demonstrating clear weaknesses against both MIAs. NLS shows a slight increase in mutual information for shallow neural networks, leading to weaker resistance to both types of attacks. Under IB-based defenses, usability sees a slight decline, but all methods except MID provide strong resistance against MLE-based MIA, with MSE values exceeding 0.5. Among these techniques, only BiDO significantly impacts MI and $\delta(z)$ , providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) are challenging, limiting its generalizability—a limitation further evidenced in subsequent experiments. This experiment adheres to the InfoSCISSORS setup, providing CLUB with a complete dataset and sufficient training iterations to ensure convergence. While CLUB effectively defends against MLE-based MIA, it struggles under the experimental conditions to resist Gen-based MIA and results in a 4.83% accuracy loss. Under AE-based defenses, target model usability improves, likely because the original model lacked sufficient capacity for optimal task classification, and the added model complexity enhanced performance. Additionally, MI indicates notable reductions in mutual information, showcasing reasonable defense effectiveness. However, in gray-box scenarios, AE-based defenses show reduced resistance, and the added complexity imposes a eightfold storage burden on edge models, as revealed by $|\theta_{\text{edge}}|$. SiftFunnel achieves the best defense performance against both MIAs while maintaining an accuracy loss below 3%. For MLE-based MIA, it achieves an MSE of 0.6792, and for Gen-based MIA, an MSE above 0.1. In gray-box scenarios, it maintains an MSE of 0.06 and an SSIM of 0.7647, ensuring robust resistance. Furthermore, MI and $\delta(z)$ demonstrate that SiftFunnel exerts optimal influence on the key elements of $D_{\text{mia}}$. Additionally, $|\theta_{\text{edge}}|$ reveals that SiftFunnel reduces the edge model’s load by nearly 20 times.
2) The effect of model architecture. Changing the model architecture does not affect the parameter settings of the proposed method, which remain consistent with Section 6.1. We adapt the model to a ResNet-18 with skip connections, using the first residual block as the edge model. This setup is challenging for defense techniques, as shallow networks are simple and do not alter spatial dimensions. Table 4 shows minimal MI impact in unprotected models, leading to weak resistance against MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$. Under these conditions, existing perturbation-based defenses fail to balance usability and privacy effectively, resulting in poor defense performance. Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by leveraging the estimation of $I(x,z)$ as a loss to optimize the edge model. However, its effectiveness remains limited, particularly against Gen-based MIA. For BiDO, despite following the kernel bandwidth estimation methods from (Peng et al., 2022) and performing extensive experimental tuning, it fails to provide a configuration that balances usability and privacy, ultimately leading to suboptimal defense outcomes. InfoSCISSORS demonstrates some impact on MI under this configuration, but the increase in $\delta(z)$ undermines its defense effectiveness against Gen-based MIA, showing only marginal improvements. AE-based defenses exhibit certain resistance to attacks on low-dimensional features in earlier stages, but under the ResNet configuration, their performance diminishes. While this method reduces $\delta(z)$ by 60%, it has minimal impact on MI, with only a 0.04% decrease. Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
@@
In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, achieving the best resistance against MIAs while maintaining usability and deployability. Notably, it achieves a minimum SSIM of 0.0793, highlighting its superior performance in balancing privacy and usability.
@@
3) The effect of data type. Changes in data types lead to shifts in task requirements. For instance, in this study, the facial recognition task involves 530 classes, while ChestX-ray classification involves 4 classes, necessitating parameter adjustments as shown in Table 5. As analyzed in the previous section on model architectures, to achieve better defense performance, we adjusted the edge model divisions for ResNet-18 and VGG. Specifically, the division points were set after the second residual block for ResNet and after the first ReLU layer following the second max-pooling layer for VGG. The results, summarized in the table 5, show that the defense performance remains unaffected by changes in data types. SiftFunnel continues to be the best defense method, offering an optimal balance between usability, privacy, and deployability.
@@
4) Ablation Study. To validate the effectiveness of SiftFunnel’s model structure and loss function design, as well as its defense capability, we conducted ablation experiments using Gen-based MIA. As shown in Table 6, we separately removed the Information Funnel and Attention Blocks from the model structure. The results indicate that removing the Funnel had minimal impact on accuracy, with only a 0.1% decrease, while removing the attention mechanism led to a significant increase in MI compared to SiftFunnel. For the loss function, we evaluated the impact of individually removing KL divergence, linear and nonlinear constraints, and the $l_{1}$-norm. The results demonstrate that KL divergence and $l_{1}$-norm moderately improved defense, whereas nonlinear constraints played a critical role; removing them increased the MSE to 0.0268. Additionally, we verified the hypothesis from Section 5.2, showing that even when adversaries reconstructed input using features from the cloud’s compensation module, SiftFunnel outperformed existing defenses in protecting transmitted features.
@@
7. Discussion
Systematic experiments validated both the criterion and the effectiveness of the method. However, our analysis revealed that achieving optimal resistance against attacks remains challenging for current defense methods, including SiftFunnel, particularly when applied to shallow edge neural networks. When edge models perform only basic feature processing, even SiftFunnel can still allow Gen-based MIA to reconstruct blurred feature structures, while other methods are more vulnerable to input reconstruction. Enhancing privacy protection for extremely shallow edge neural networks presents a promising direction for future research.
Additionally, using MI as an evaluation metric, we observed that some defense methods result in only a 0.5% reduction in MI between inputs and features, yet current MLE-based MIAs still fail to bypass these defenses. This is directly linked to the lack of advancement in existing attack methods. Based on our experimental findings, we suggest that further exploration of more powerful MIAs in white-box scenarios could be a potential research direction.
@@ 8. Conclusion
The increasing complexity of neural networks and inference tasks, alongside the demand for efficiency and real-time feedback, presents significant challenges for edge devices with limited resources. Collaborative inference offers a solution by delegating feature extraction to edge devices and offloading subsequent tasks to the cloud, but this process leaves transmitted features vulnerable to MIAs. To address these vulnerabilities, we established a first criterion $D_{\text{mia}}$ for evaluating MIA difficulty and proposed SiftFunnel, a privacy-preserving framework designed to limit redundant information and enhance usability. By integrating linear and non-linear correlation constraints, label smoothing, and a funnel-shaped edge model with attention mechanisms, SiftFunnel achieves robust privacy protection while maintaining usability and deployability. Experimental results demonstrate its effectiveness, achieving superior defense against MIAs with minimal accuracy loss and a balanced trade-off among usability, privacy, and practicality.

### v4

@@ 6.2. Evaluation Metrics
6.3. Experimental Results
The results are analyzed from 4 perspectives: the impact of defense methods on $I(x,z)$ and $\delta(z)$, the influence of datasets and model architectures, and ablation studies.
@@ 6.3. Experimental Results
1) The effect of various defenses on the MI and $\delta(z)$. As shown in Table 2, defense effectiveness is reflected by MI and $\delta(z)$, aligning with the criterion $D_{\text{mia}}$. Specifically, unprotected edge models are highly vulnerable to MIAs, with MI at 0.0566, MSE below 0.01, and SSIM exceeding 0.98. Perturbation-based defenses reduce MI by at most 11% but cause over a 10% drop in Test ACC, with NLS even increasing MI in shallow networks, weakening resistance to attacks. UIB-based defenses slightly reduce usability but effectively counter MLE-based MIAs, except for MID, which shows weaker resistance. Among them, BiDO significantly impacts MI and $\delta(z)$, providing effective defense, but its parameter selection and kernel width configuration (as shown in the table) limits generalizability, as confirmed in subsequent experiments. This experiment follows the InfoSCISSORS setup, ensuring CLUB receives a complete dataset and sufficient training for convergence. While CLUB effectively defends against MLE-based MIAs, it struggles against Gen-based MIAs and incurs a 4.83% accuracy loss. AE-based defenses improve model usability, likely due to enhanced model capacity, while MI reductions confirm reasonable defense effectiveness. However, gray-box scenarios weaken AE-based defenses, and their added complexity increases storage costs by 8×8$\times$8 ×, as seen in $|\theta_{\text{edge}}|$. SiftFunnel achieves the best defense performance against both MIAs while keeping accuracy loss below 3%. It achieves MSE of 0.6792 (MLE-based MIA) and above 0.1 (Gen-based MIA). In gray-box scenarios, it maintains MSE of 0.06 and SSIM of 0.7647, ensuring robust resistance. Furthermore, MI and $\delta(z)$ confirm its optimal impact on the key factors of $D_{\text{mia}}$. Additionally, $|\theta_{\text{edge}}|$ reveals a nearly 20×20$\times$20 × reduction in edge model load, highlighting SiftFunnel’s efficiency.
@@
2) The effect of model architecture. Modifying the model architecture does not alter the parameter settings of the proposed method, which remain consistent with Section 6.1. We adapt the model to ResNet-18 with skip connections, using the first residual block as the edge model. This setup poses challenges for defense techniques, as shallow networks retain simple structures and do not modify spatial dimensions. Table 3 shows that unprotected models exhibit minimal MI impact, resulting in weak resistance to MIAs, with MLE-based MIA achieving an MSE of $10^{-4}$ and Gen-based MIA $10^{-3}$. Under these conditions, perturbation-based defenses fail to balance usability and privacy, leading to poor defense performance. Among IB-based defenses, only VIB demonstrates resistance to MLE-based MIA by optimizing the edge model through $I(x,z)$ estimation as a loss function, but its effectiveness remains limited, particularly against Gen-based MIA. For BiDO, despite applying kernel bandwidth estimation methods from (Peng et al., 2022) and extensive experimental tuning, it fails to achieve a configuration that effectively balances usability and privacy, resulting in suboptimal defense performance. InfoSCISSORS impacts MI under this configuration, but the increase in $\delta(z)$ undermines its effectiveness against Gen-based MIAs, leading to only marginal improvements. AE-based defenses show some resistance to attacks on low-dimensional features in earlier stages, but their effectiveness diminishes under the ResNet configuration. While they reduce $\delta(z)$ by 60%, they have minimal impact on MI, achieving only a 0.04% reduction. Consequently, AE-based defenses struggle to counter both black-box and gray-box Gen-based MIAs effectively.
Table 6. Quantitative evaluation of the resistance to MIA across different architectures on CIFAR-10. The edge model divisions for CNN and ResNet-18 follow the configurations in Tables 2 and 3, while for VGG16, the edge model is set at the first ReLU layer after a max-pooling operation.
MLE-based Attack (He et al., 2019) Gen-based Attack (He et al., 2020)
0.1166 14.1112 0.5419
0.1957 11.8562 0.0793
0.0827 15.5928 0.6559
@@
In contrast, SiftFunnel reduces MI by 50% and lowers $\delta(z)$ to the millesimal scale, delivering the strongest resistance against MIAs while preserving usability and deployability. Notably, it achieves a minimum SSIM of 0.0793, underscoring its superior ability to balance privacy and usability. The different protection performances of the three model architectures with the same dataset are shown in Table 6.
@@
3) The effect of data type. Changes in data types lead to shifts in task requirements. For instance, in this study, the facial recognition task involves 530 classes, while ChestX-ray classification involves 4 classes, necessitating parameter adjustments as shown in Table 4. As analyzed in the previous section on model architectures, to achieve better defense performance, we adjusted the edge model divisions for ResNet-18 and VGG. Specifically, the division points were set after the second residual block for ResNet and after the first ReLU layer following the second max-pooling layer for VGG. The results, summarized in Table 4 and Figure 5, show that the defense performance remains unaffected by changes in data types. SiftFunnel continues to be the best defense method, offering an optimal balance between usability, privacy, and deployability.
As shown in Figure 5, the VGG16 model protected by SiftFunnel demonstrates better resistance to Gen-based MIA in terms of the visual effects of usability and privacy on facial and X-ray data, with PSNR values consistently around 15.
@@
4) Ablation Study. To assess the effectiveness of SiftFunnel’s model structure, loss function design, and defense capability, we conducted ablation experiments using Gen-based MIA. As shown in Table 5, we separately removed the Information Funnel and Attention Blocks from the model structure. The results indicate that removing the Funnel had minimal impact on accuracy, reducing it by only 0.1%, whereas removing the attention mechanism significantly increased MI, weakening privacy protection. For the loss function, we examined the effect of removing KL divergence, linear and nonlinear constraints, and the $l_{1}$-norm. The results show that KL divergence and $l_{1}$-norm provided moderate defense improvements, while nonlinear constraints were critical—removing them increased MSE to 0.0268. Additionally, we validated the hypothesis from Section 5.2, demonstrating that even when adversaries reconstructed inputs using cloud-compensated features, SiftFunnel still outperformed existing defenses in protecting transmitted features.
Figure 5. Quantitative and visual evaluation of privacy and usability against Gen-based MIA in the VGG16 edge model on the ChestX-ray ($\lambda_{1}=3.5,\lambda_{2}=1.2,\lambda_{3}=0.8$) and FaceScrub ($\lambda_{1}=4.5,\lambda_{2}=0.4,\lambda_{3}=0.1$).
[…]
5
@@ 7. Conclusion
While CI enhances efficiency, it also exposes transmitted features to MIAs, posing severe privacy risks. Despite extensive research, a fundamental understanding of the factors determining MIA effectiveness remains lacking, and existing defenses struggle to balance usability, privacy, and deployability, facing theoretical, methodological, and practical constraints. To address these challenges, we proposed the first criterion for evaluating MIA difficulty in CI, identifying MI, entropy, and effective information volume as key influencing factors. Based on this, we proposed SiftFunnel, a privacy-preserving framework designed to limit redundant information while maintaining usability and deployability. By integrating linear and nonlinear correlation constraints, label smoothing, and a funnel-shaped edge model with attention mechanisms, SiftFunnel achieved robust privacy protection with an average accuracy loss of around 3%, and revealed a nearly 20×20$\times$20 × reduction in edge burdens. Experimental results demonstrated that SiftFunnel offers superior privacy protection across various evaluation metrics and achieves an optimal balance among usability, privacy, and practicality.